OpenStudy (anonymous):

what's the derivative of e^(sin(1)x)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[e^{\sin(1)x}\]??

OpenStudy (anonymous):

its the same thing cause of the exponential rule

OpenStudy (anonymous):

if that is really the question, \[\sin(1)\] is a number so your answer is \[\sin(1)e^{\sin(1)x}\]

OpenStudy (agreene):

if you are asking: \[\frac{d}{dx}e^{sinx}\] then: \[e^{sinx}cosx\]